OpenStudy (anonymous):

An equation and the steps to solve it are shown below: 2x - one over four of (12x + 4) = 3 Step 1: 2x + 3x + 1 = 3 Step 2: 5x + 1 = 3 Step 3: 5x + 1 - 1 = 3 - 1 Step 4: 5x = 2 In which step was the error made and why? Step 1, because incorrect signs have been used Step 2, because the terms have been added incorrectly Step 3, because a number has been subtracted from both sides of the equation No error was made @ikram002p

OpenStudy (anonymous):

@donka74

OpenStudy (donka74):

i think its step 2 i really dont know this one

OpenStudy (anonymous):

do you know? @ikram002p

OpenStudy (ikram002p):

I think if the equation was \[\frac{ 2x-1 }{ 4(12x+4)} = 3\] then it would be wrong from step one but if the equation was \[2x + 3x + 1 = 3 \] then the anwer would be No error was made